Imagine a population that has 124 A1A1 individuals, 160 A1A2 individuals, and 27 A2A2 individuals. Male A1A1 is the preferred mate of females and gets all the matings. What would the genotype frequencies be for the A1A2 offspring of this population? Answer to 2 decimal places.  
Answer #1

Representation:

  • A1A1= Homozygous dominant =124
  • A2A2 = Homozygous recessive = 27
  • A1A2= Heterozygous = 160
  • Total population: 124+27+160=311
  • Frequency for A1 allele = (124+{1/2 x160})/ 311 =0.66 (approx)
  • Frequency for A2 allele = (27+{1/2x 160})/ 311 =0.34 (approx)
  • Thus, cross of A1A1 x A2A2 produce A1A2
  • Probability = 0.66 x 0.34 = 0.22 (approx).